# bibCyclic.bib

### v2

@@ 5.2 Removing $l$-strips
Here we let $\mathsf{v}$ and $\mathsf{w}$ be arbitrary. Let $x\in\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{T}}$. Recall from Proposition 3.2 that $x$ is equivalent to a certain tuple of partitions. Assume that not all of these partitions are $l$-core partitions. So there is some partition $\lambda$ which has a removable $l$-strip $\alpha$. Let $\lambda^{\prime}$ be a partition obtained by removing $\alpha$ from $\lambda$, i.e., $\lambda^{\prime}=\lambda\setminus\alpha$. The tuple of partitions for $x$, but with $\lambda$ replaced by $\lambda^{\prime}$, indexes a $\mathsf{T}$-fixed point $x^{\prime}$ on some quiver variety $\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w})^{\mathsf{T}}$.
Let $f\in\text{QM}_{\text{ns }0}\left(\mathbb{P}^{1}\to\mathcal{M}(\mathsf{v},\mathsf{w})\right)$ be a quasimap such that $f(0)=x$. Recall, see for example [dinkinsthesis] section 5.1.3, that $f$ is equivalent to a tuple of reverse plane partitions over the partitions for $x$. By disregarding the boxes over $\alpha$ in this tuple of reverse plane partitions, we obtain $f^{\prime}\in\text{QM}_{\text{ns }0}\left(\mathbb{P}^{1}\to\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w})\right)$ such that $f^{\prime}(0)=x^{\prime}$.
The quasimaps $f$ and $f^{\prime}$ provide vector bundles $\mathscr{V}_{i,f}$ and $\mathscr{V}_{i,f^{\prime}}$ which, by construction, are related by the following lemma.
@@
$-\sum_{i}\left((\mathsf{v}_{i}-1)d_{i}-\deg\mathscr{V}_{i}^{\prime}\right)$
@@
$\sum_{i}\left((\mathsf{v}_{i-1}-1)d_{i}-\deg\mathscr{V}_{i-1}^{\prime}+\mathsf{w}_{i}d_{i}\right)$
@@
As the base case, we consider the situation where there is a single $l$-strip among the two partitions. For concreteness, we assume $\lambda_{0}=\emptyset$ and $\lambda_{m}$ is a single $l$-strip, which in this case must be a hook with $l$ boxes. A similar argument handles the case where $\lambda_{0}$ is an $l$-strip and $\lambda_{m}=\emptyset$. Let $f\in\left(\text{QM}_{\text{ns }0}\right)^{\mathbb{C}^{\times}_{q}}$ be a nonconstant quasimap such that $f(0)=x$. Write $\mathscr{V}_{i}=\mathcal{O}(d_{i})$. Let $d^{\prime}$ and $d^{\prime\prime}$ the degrees corresponding to the two outermost boxes in the hook. Since $f$ is nonconstant, at least one of $d^{\prime}$ or $d^{\prime\prime}$ must be strictly positive. Also, $\deg\mathscr{T}^{1/2}_{p,f}$, having no negative terms, is at least $d_{0}+d_{m}+(d^{\prime\prime}-d_{m})+\max(0,d^{\prime}-d^{\prime\prime})$. So
$\deg\mathscr{V}_{m,f}-\deg\mathscr{T}^{1/2}_{p,f}\leq d_{m}-d_{0}-d^{\prime\prime}-\max(0,d^{\prime}-d^{\prime\prime})\\ <-d^{\prime\prime}-\max(0,d^{\prime}-d^{\prime\prime})<0$
@@
For the inductive step, we let $\lambda_{0}$ and $\lambda_{m}$ be arbitrary partitions, at least one of which contains a removable $l$-strip. Let $f$ be a nonconstant $\mathbb{C}^{\times}_{q}$-fixed quasimap such that $f(0)=x$. As in the first paragraph of this subsection, we can remove an $l$-strip to obtain a fixed point $x^{\prime}$ on another quiver variety, along with a quasimap $f^{\prime}$ with $f^{\prime}(0)=x^{\prime}$.
@@
Finally, suppose that $x^{\prime}$ and $f^{\prime}$ cannot be chosen in such a way that $f^{\prime}$ is nonconstant. This implies that only one of $\lambda_{0}$ or $\lambda_{m}$ contains a removable $l$-strip and that the other is empty. Without loss of generality, we can assume that $\lambda_{0}=\emptyset$. Furthermore, all the boxes in the reverse plane partition over $\lambda_{m}$ corresponding to $f$ lie over a single $l$-strip $\alpha$ of $\lambda_{m}$. For two boxes $a,b\in\alpha$, we write $a\to b$ if $a$ is an inner corner, $b$ is an outer corner, $b$ lies in the hook based at $a$, and $c(b)>c(a)$. Let $d_{l}$ and $d_{r}$ be the degrees corresponding to the boxes of $\alpha$ with minimal and maximal content, respectively. Let $d_{a}$ be the degree for a box $a\in\alpha$, and let $d_{0}$ and $d_{m}$ be the degrees for the (unique) boxes in $\alpha$ of content 00 and $m$. Then
@@ 6 Fusion operator and quantum toroidal algebra
In [OkSm], a certain quantum group was constructed which acts on the equivariant $K$-theory of quiver varieties. The main result of [OkSm] identifies the quantum difference operator with an operator constructed using this quantum group. In [zhu2], the quantum group was identified with quantum toroidal $\mathfrak{sl}_{n}$. So we can apply known results about this algebra to the enumerative geometry of quiver varieties. Our goal in this section is Proposition 6.8, and we introduce the minimum amount needed to get there.
@@ 6.1 Quantum toroidal algebra
We recall a few basic facts about $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$, the quantum toroidal algebra of $\mathfrak{g}=\mathfrak{sl}_{l}$, following [Tsym] and [Wen]. Inside $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$, there are two copies of the quantum affine algebra $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$ which are denoted by $h,v:\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})\hookrightarrow\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ and referred to as the “horizontal” and “vertical” subalgebras.
Let
$\mathsf{F}:=\bigoplus_{\mathsf{v}\in\mathbb{Z}_{\geq 0}^{l}}K_{\mathsf{T}}(\mathcal{M}(\mathsf{v},\delta_{0}))_{loc}$
The quantum group of [OkSm] is defined as follows. Using stable envelopes, [OkSm] defines a collection of $R$-matrices acting on $\mathsf{F}\otimes\mathsf{F}$. Using the FRT procedure of [FRT], their quantum group is defined as the algebra generated by all matrix coefficients of these $R$-matrices. This quantum group is generated by matrix coefficients of the so-called wall $R$-matrices which we will denote by $\mathcal{R}^{\text{OS}}_{w}$ for $w\in\mathbb{Q}^{l}$.
The algebra $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ also acts on $\mathsf{F}$. We denote this representation by $\rho_{f}:\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})\to\text{End}(\mathsf{F})$ since it is known as the fermionic Fock representation. In particular, the $R$-matrix of the horizontal subalgebra acts on $\mathsf{F}\otimes\mathsf{F}$. We write $\mathcal{R}$ for the universal quasi $R$-matrix of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$.
Theorem 6.1 ([zhu2] Theorem 7.2).
$\mathcal{R}^{\text{OS}}_{0}=(\rho_{f}\otimes\rho_{f})(h(\mathcal{R}))$.
@@
To be clear, [zhu2] proves a much stronger result. Nevertheless, we only require information about the $R$ matrix for wall 00.
There is another representation of $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ known as the bosonic Fock representation: $\rho_{b}:\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})\to\text{End}\left(\Lambda_{t_{1},t_{2}}^{(l)}\otimes_{\mathbb{Q}}\mathbb{Q}[Q]\right)$, see [saito] and [Wen] for the construction of the representation. Although we have an isomorphism of vector spaces from Theorem 3.4, $\mathsf{F}\cong\Lambda_{t_{1},t_{2}}^{(l)}\otimes_{\mathbb{Q}}\mathbb{Q}[Q]$, the bosonic and fermionic representations differ under this identification by a twist known as the Miki automorphism.
@@ 6.2 Khoroshkin-Tolstoy formula
By the Khoroshkin-Tolstoy formula, see [KT, KT2] and also [zhu1] section 2.8, the quasi $R$-matrix of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$ can be written in the form
@@
Recall that, among the generators of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$, there are the coefficients of the currents $e_{i}(z)=\sum_{k\in\mathbb{Z}}e_{i,k}z^{-k}$ and $f_{i}(z)=\sum_{k\in\mathbb{Z}}f_{i,k}z^{-k}$ for $11\leq i\leq l-$ 1. Recall from [Wen] that $\mathcal{R}^{+}=1+\ldots$ (resp. $\mathcal{R}^{-}=1+\ldots$) where the dots stand for an operator whose first (resp. second) tensorands are products of the coefficients of the $e$-currents and whose second (resp. first) tensorands are products of the coefficients of the $f$-currents.
We need the following formula for the action of $h(\mathcal{R}_{0})$ in the tensor product of two Fock representations.
@@
$(\rho_{f}\otimes\rho_{f})h(\mathcal{R}_{0})=\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}(\hbar^{n/2}-\hbar^{-n/2})p^{(i)}_{n}\otimes\frac{\partial}{\partial p^{(i)}_{n}}\right)$
@@
We follow the notation of [Wen], up to a transposition of tensor factors. By Proposition 3.4 of [Wen],
@@
  • $\omega$ stands for the Miki automorphism of the quantum toroidal algebra, for which it is known that $\omega\circ h=v$ and, under the identification of Proposition 3.4, $\rho_{b}\circ\omega=\rho_{f}$, see [Miki1, Miki2, Wen].
@@
$(\rho_{f}\otimes\rho_{f})h(\mathcal{R}_{0})$ $=(\rho_{b}\otimes\rho_{b})\circ(\omega\otimes\omega)h(\mathcal{R}_{0})$

### v3

@@ 5.3 Removing $l$-strips
Let $\mathsf{v}$ and $\mathsf{w}$ be arbitrary. Let $x\in\mathcal{M}(\mathsf{v},\mathsf{w})^{\mathsf{T}}$. Recall from Proposition 3.2 that $x$ is equivalent to a certain tuple of partitions. Assume that not all of these partitions are $l$-core partitions. So there is some partition $\lambda$ which has a removable $l$-strip $\alpha$. Let $\lambda^{\prime}$ be a partition obtained by removing $\alpha$ from $\lambda$, i.e., $\lambda^{\prime}=\lambda\setminus\alpha$. The tuple of partitions for $x$, but with $\lambda$ replaced by $\lambda^{\prime}$, indexes a $\mathsf{T}$-fixed point $x^{\prime}$ on some quiver variety $\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w})^{\mathsf{T}}$.
Let $f\in\text{QM}_{\text{ns }0}\left(\mathbb{P}^{1}\to\mathcal{M}(\mathsf{v},\mathsf{w})\right)^{\mathsf{T}\times\mathbb{C}^{*}_{q}}$ be a quasimap such that $f(0)=x$. As we recalled above, $f$ is equivalent to a tuple of reverse plane partitions over the partitions for $x$. By disregarding the boxes over $\alpha$ in this tuple of reverse plane partitions, we obtain $f^{\prime}\in\text{QM}_{\text{ns }0}\left(\mathbb{P}^{1}\to\mathcal{M}(\mathsf{v}^{\prime},\mathsf{w})\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$ such that $f^{\prime}(0)=x^{\prime}$.
The quasimaps $f$ and $f^{\prime}$ provide vector bundles $\mathscr{V}_{i,f}$ and $\mathscr{V}_{i,f^{\prime}}$. By construction, $\mathscr{V}_{i,f}=\mathscr{V}_{i,f^{\prime}}\oplus\mathcal{O}(d_{i})$ where $d_{i}\geq 0$ is the component of the reverse plane partition for $f$ corresponding to the unique box in $\alpha$ of content $i$. We remind the reader of our convention for contents of boxes in tuples of partitions given in section 3.3. The following is clear.
@@
$\sum_{i}\deg\text{Hom}\left(\mathscr{V}_{i}^{\prime},\mathcal{O}(d_{i})\right)=-\sum_{i}\left((\mathsf{v}_{i}-1)d_{i}-\deg\mathscr{V}_{i}^{\prime}\right)$
@@
$\sum_{i}\deg\text{Hom}(\mathscr{V}_{i-1}^{\prime},\mathcal{O}(d_{i}))+\deg\text{Hom}(\mathcal{O}(0)^{\oplus\mathsf{w}_{i}},\mathcal{O}(d_{i}))\\ =\sum_{i}\left((\mathsf{v}_{i-1}-1)d_{i}-\deg\mathscr{V}_{i-1}^{\prime}+\mathsf{w}_{i}d_{i}\right)$
@@
As the base case, we consider the situation where there is a single $l$-strip among the two partitions. For concreteness, we assume $\lambda_{0}=\emptyset$ and $\lambda_{m}$ is a single $l$-strip, which in this case must be a hook with $l$ boxes. A similar argument handles the case where $\lambda_{0}$ is an $l$-strip and $\lambda_{m}=\emptyset$. Let $f\in\left(\text{QM}_{\text{ns }0}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$ be a nonconstant quasimap such that $f(0)=x$. Write $\mathscr{V}_{i}=\mathcal{O}(d_{i})$. Let $d^{\prime}$ and $d^{\prime\prime}$ the degrees corresponding to the two outermost boxes in the hook. Since $f$ is nonconstant, at least one of $d^{\prime}$ or $d^{\prime\prime}$ must be strictly positive. Also,
$\deg\mathscr{T}^{1/2}_{p,f}=d_{0}+d_{m}+\sum_{\begin{subarray}{c}i\\ d_{i+1}>d_{i}\end{subarray}}(d_{i+1}-d_{i})=d_{0}+d_{m}+d^{\prime\prime}-d_{m}+\max(0,d^{\prime}-d^{\prime\prime})$
where the first equality is by definition and the second is because $\lambda_{m}$ is a hook. Since $\deg\mathscr{V}_{m,f}=d_{m}$, we have
$\deg\mathscr{V}_{m,f}-\deg\mathscr{T}^{1/2}_{p,f}\leq d_{m}-d_{0}-d^{\prime\prime}-\max(0,d^{\prime}-d^{\prime\prime})\\ \leq-d^{\prime\prime}-\max(0,d^{\prime}-d^{\prime\prime})<0$
@@
For the inductive step, we let $\lambda_{0}$ and $\lambda_{m}$ be arbitrary partitions, at least one of which contains a removable $l$-strip. Let $f$ be a nonconstant $\mathsf{T}\times\mathbb{C}^{\times}_{q}$-fixed quasimap such that $f(0)=x$. As in the first paragraph of this subsection, we can remove an $l$-strip to obtain a fixed point $x^{\prime}$ on another quiver variety, along with a quasimap $f^{\prime}$ with $f^{\prime}(0)=x^{\prime}$.
@@
Finally, suppose that $x^{\prime}$ and $f^{\prime}$ cannot be chosen in such a way that $f^{\prime}$ is nonconstant. This implies that only one of $\lambda_{0}$ or $\lambda_{m}$ contains a removable $l$-strip and that the other is empty. Without loss of generality, we can assume that $\lambda_{0}=\emptyset$. Furthermore, all the boxes in the reverse plane partition over $\lambda_{m}$ corresponding to $f$ lie over a single $l$-strip $\alpha$ of $\lambda_{m}$. For two boxes $a,b\in\alpha$, we write $a\to b$ if $a$ is an inner corner, $b$ is an outer corner, $b$ lies in the hook based at $a$, and $c(b)>c(a)$. Let $d_{l}$ and $d_{r}$ be the degrees corresponding to the leftmost and rightmost boxes of $\alpha$, respectively. Let $d_{a}$ be the degree for a box $a\in\alpha$, and let $d_{0}$ and $d_{m}$ be the degrees for the (unique) boxes in $\alpha$ of content 0 and $m$. Then
@@ 6 Fusion operator and quantum toroidal algebra
In [48], Okounkov and Smirnov constructed a certain quantum group $\mathscr{U}^{\text{OS}}$ which acts on the equivariant $K$-theory of quiver varieties. The main result of [48] identifies the quantum difference operator with an operator inside (a certain completion) of $\mathscr{U}^{\text{OS}}$. In [68], it was proven that $\mathscr{U}^{\text{OS}}\cong\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$, where $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ denotes the quantum toroidal algebra for $\mathfrak{g}=\mathfrak{sl}_{l}$. So we can apply known results about $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ to the enumerative geometry of quiver varieties. Our goal in this section is Proposition 6.10, and we introduce the minimum amount needed to get there.
@@ 6.1 Quantum toroidal algebra
We recall a few basic facts about $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$, the quantum toroidal algebra of $\mathfrak{g}=\mathfrak{sl}_{l}$, following [63] and [66]. Inside $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$, there are two copies of the quantum affine algebra $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$ which are denoted by $h,v:\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})\hookrightarrow\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ and referred to as the “horizontal” and “vertical” subalgebras.
There are three representations of $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ that we will consider. The first is the “geometric representation”
$\rho_{g}:\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})\to\text{End}(\mathsf{G}),\quad\mathsf{G}:=\bigoplus_{\mathsf{v}\in\mathbb{Z}_{\geq 0}^{l}}K_{\mathsf{T}}(\mathcal{M}(\mathsf{v},\delta_{0}))_{loc}$
which was constructed in [57], see also [44]. The $K$-theory classes of skyscraper sheaves $I_{\lambda}]$ of fixed points give a basis of this representation.
The quantum group $\mathscr{U}^{\text{OS}}$ of [48] is defined as follows. Using stable envelopes, [48] defines a collection of $R$-matrices acting on $\mathsf{G}\otimes\mathsf{G}$. Using the FRT procedure of [19], $\mathscr{U}^{\text{OS}}$ is defined as the algebra generated by all matrix coefficients of these $R$-matrices. It is known that $\mathscr{U}^{\text{OS}}$ is generated by matrix coefficients of the so-called wall $R$-matrices which we will denote by $\mathcal{R}^{\text{OS}}_{w}$ for $w\in\mathbb{Q}^{l}$.
In particular, the $R$-matrix of the horizontal subalgebra acts on $\mathsf{G}\otimes\mathsf{G}$. We write $\mathcal{R}$ for the universal quasi $R$-matrix of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$.
Theorem 6.1 ([68] Theorem 7.2).
$\mathcal{R}^{\text{OS}}_{0}=(\rho_{g}\otimes\rho_{g})(h(\mathcal{R}))$.
@@
To be clear, [68] proves a much stronger result. Nevertheless, we only require information about the $R$ matrix for wall 0.
The second $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ representation that we consider is the “fermionic Fock” representation $\rho_{f}:\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})\to\text{End}(\mathsf{F})$, where $\mathsf{F}$ is the vector space over $\mathbb{Q}(t_{1},t_{2})$ with basis given by partitions $|\lambda\rangle$. For the definition of the $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ action, see [66] section 3.5.
Third, there is the “bosonic” representation [55] (also called the vertex representation)
$\rho_{b}:\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})\to\text{End}(\mathsf{B}),\quad\mathsf{B}:=\Lambda_{t_{1},t_{2}}^{(l)}\otimes_{\mathbb{Q}}\mathbb{Q}[Q]$
which has a basis given by wreath Macdonald polynomials $H_{\lambda}$.
Tsymbaliuk proved that there is an isomorphism $T:\mathsf{F}\xrightarrow{\sim}\mathsf{B}$ which intertwines $\rho_{f}$ with $\rho_{b}\circ\omega$, where $\omega$ is the so-called Miki automorphism of $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ [63]. Wen proved that $T(|\lambda\rangle)=a_{\lambda}H_{\lambda}$ for some $a_{\lambda}\in\mathbb{Q}(t_{1},t_{2})$ [66]. The scalars $a_{\lambda}$ are difficult to compute, and a formula for them is not known [65].
On the other hand, there exists $b_{\lambda}\in\mathbb{Q}(t_{1},t_{2})$ such that the map $\mathsf{G}\to\mathsf{F}$ induced by $I_{\lambda}]\mapsto b_{\lambda}|\lambda\rangle$ is an isomorphism of $\mathscr{U}_{\hbar}(\ddot{\mathfrak{g}})$ representations, see [39]. Combinatorial formulas for $b_{\lambda}$ are known.
Conjecture 6.3.
We have the equality $a_{\lambda}=b_{\lambda}^{-1}$.
Assumption: We henceforth assume Conjecture 6.3 holds. This implies that the isomorphism $\mathsf{G}\xrightarrow{\sim}\mathsf{B}$ induced by $I_{\lambda}]\to H_{\lambda}$ intertwines $\rho_{g}$ with $\rho_{b}\circ\omega$.
@@ 6.2 Khoroshkin-Tolstoy formula
By the Khoroshkin-Tolstoy formula, see [29, 30] and also [67] section 2.8, the quasi $R$-matrix of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$ can be written in the form
@@
This factorization depends on a choice of normal ordering of the positive affine roots. Let $\Delta$ denote the affine roots. We denote the affine simple roots by $\alpha_{0},\ldots,\alpha_{l-1}$ and the imaginary root by $\delta$. Let $\Delta_{+}^{\text{re}}$ denote the real positive roots. Let $\Delta^{\text{fin}}$ be the finite root system with simple roots $\alpha_{1},\ldots,\alpha_{l-1}$, and let $\Delta^{\text{fin}}_{+}$ denote the positive finite roots. We choose a normal ordering, denoted $\leqslant$, such that all of the positive roots of the finite root system precede the imaginary root.
Then [30] constructs a quantum Cartan-Weyl basis for $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$. It consists of elements $e_{\gamma}\in\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$ for each $\gamma\in\Delta$. In terms of this basis, the factors $\mathcal{R}^{\pm}$ of the $R$-matrix are
$\mathcal{R}^{+}=\prod_{\begin{subarray}{c}\gamma\in\Delta_{+}^{\text{re}}\\ \gamma<\delta\end{subarray}}\exp_{\hbar^{-(\gamma,\gamma)}}\left(a(\gamma)e_{\gamma}\otimes e_{-\gamma}\right),\qquad a(\gamma)\in\mathbb{C}(\hbar)$ (12)
and
$\mathcal{R}^{-}=\prod_{\begin{subarray}{c}\gamma\in\Delta_{+}^{\text{re}}\\ \gamma>\delta\end{subarray}}\exp_{\hbar^{-(\gamma,\gamma)}}\left(a(\gamma)e_{\gamma}\otimes e_{-\gamma}\right),\qquad a(\gamma)\in\mathbb{C}(\hbar)$ (13)
where the products are ordered according to the choice of normal ordering and $\exp$ denotes
$\exp_{b}(a)=\sum_{n\geq 0}\frac{a^{n}}{(n)_{q}},\quad(n)_{b}:=\frac{1-b^{n}}{1-b}$
By our assumption on the normal ordering, the roots $\gamma$ appearing in the product for $\mathcal{R}^{+}$ are of the form $\alpha+k\delta$ where $\alpha\in\Delta^{\text{fin}}_{+}$ and $k\geq 0$. Similarly, all the roots appearing in the product for $\mathcal{R}^{-}$ are of the form $\alpha_{0}+\alpha+k\delta$ where $\alpha\in\Delta^{\text{fin}}_{+}$ and $k\geq 0$.
@@
Let $\gamma\in\Delta^{\text{re}}_{+}$ and write $\gamma=\beta+k\delta$ where $k\geq 0$ and either $\beta\in\Delta^{\text{fin}}_{+}$ or $\beta=\alpha_{0}+\alpha$ for some $\alpha\in\Delta^{\text{fin}}_{+}$. Let $\operatorname{supp}(\beta)\subset\{0,\ldots,l-1\}$ be the indices of the simple roots appearing in $\beta$ with nonzero coefficient. Then $\rho_{g}(h(e_{\gamma}))$ (resp. $\rho_{g}(h(e_{-\gamma}))$) acts on $\mathsf{G}$ in the basis $I_{\lambda}]$ by adding (resp. removing) boxes of $\lambda$ with content in $\operatorname{supp}(\gamma)$.
@@
It is proven in [30] that elements of the quantum Cartan-Weyl basis of the form $e_{\alpha_{i}+k\delta}$ (resp. $e_{-(\alpha_{i}+k\delta)}$) match, up to constants, the coefficients of the $i$th generating current in the positive (resp. negative) half of Drinfeld’s realization of $\mathscr{U}_{\hbar}(\dot{\mathfrak{g}})$. From the definition of the Fock representation, see Theorem 3.18 of [66], $e_{\alpha_{i}+k\delta}$ (resp. $e_{-(\alpha_{i}+k\delta)}$) thus acts by adding (resp. removing) boxes of content $i$.
Let $\gamma=\beta+k\delta\in\Delta^{\text{re}}_{+}$ be as in the statement of the lemma. The quantum Cartan-Weyl basis is constructed inductively. In particular, if $\beta$ is not a simple root, then $e_{\beta+k\delta}$ is a $\hbar$-twisted commutator of $e_{\beta^{\prime}+k^{\prime}\delta}$ and $e_{\beta^{\prime\prime}+k^{\prime\prime}\delta}$ for some $k^{\prime},k^{\prime\prime}\geq 0$ and some roots $\beta^{\prime},\beta^{\prime\prime}$ whose supports are contained strictly in $\operatorname{supp}(\beta)$. So the claim follows from the previous paragraph by induction on the size of the support. ∎
We need the following formula for the action of $h(\mathcal{R}_{0})$ in the tensor product of two geometric representations.
Lemma 6.5.
Under the identification $\mathsf{G}\cong\mathsf{B}$, we have
$(\rho_{g}\otimes\rho_{g})h(\mathcal{R}_{0})=\exp\left(\sum_{i=0}^{l-1}\sum_{n\geq 1}(\hbar^{n/2}-\hbar^{-n/2})p^{(i)}_{n}\otimes\frac{\partial}{\partial p^{(i)}_{n}}\right)$
Proof.
We follow the notation of [66], up to a transposition of tensor factors. By Proposition 3.4 of [66],
@@
  • $\omega$ stands for the Miki automorphism of the quantum toroidal algebra, for which it is known that $\omega\circ h=v$ (see [36, 37, 66]), and by Conjecture 6.3, $\rho_{b}\circ\omega=\rho_{g}$.
@@
$(\rho_{g}\otimes\rho_{g})h(\mathcal{R}_{0})$ $=(\rho_{b}\otimes\rho_{b})\circ(\omega\otimes\omega)h(\mathcal{R}_{0})$